OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the x variable. x - 3y = 12 2x + 5y = 2

OpenStudy (kamille):

You need to multiply x-3y by -2

OpenStudy (anonymous):

a. Multiply the first equation by 2. b. Multiply the second equation by 2. c. Multiply the first equation by -2. d. Multiply the second equation by -2.

OpenStudy (anonymous):

those are the options

OpenStudy (kamille):

c

OpenStudy (anonymous):

and can you help me with some others after this

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x - 6y = -15 x - 3y = -3

OpenStudy (anonymous):

a. x - 3(6y - 15) = -3 b. x - 3(-6y - 15) = -3 c. (6y - 15) - 3y = -3 d. (-6y - 15) - 3y = -3

OpenStudy (kamille):

Well, do you know what "substitution" method means?

OpenStudy (anonymous):

yes i know

OpenStudy (jiteshmeghwal9):

First equal x variable in both equations